problem set in analytical chemistry spectroscopy

: Q. solution of fluorene (diphenylene methane) in benzene can be analyzed by making use of its absorbance at 301 m µ, where C =1.1 x 104 . if a solution of fluorene of unknown concentration in benzene exhibits an absorbance of 0.720 in a 1.00 cm cell. What is the concentration of fluorene?
Answer #1

There is relationship between absorbance and concentration named as beer Lambert law. The equation is given as:

A=ElC

A= absorbance

E= extinction coefficient

l= length of sample

C= concentration of sample

Extinction coefficient is same for same sample hence we can obtain value of E and can put in another case to obtain concentration of unknown sample

A1=301mu=301×10^(-11) cm

C1=1.1×10^4

L= 1cm

Therefore, E=[301×10^(-11)]/[1×1.1×10^4]=27.3×10^(-14)

C2=?

C2=A2/El

=0.720/[27.3×10^(-14)×1]=2.6×10^12